the time allowed for completing a certain piece of work is 40 days. By extending the time by 10 days we find that 25 fewer man will be needed. How many men are required to finish the work in 40 days?
Answer by ankor@dixie-net.com(22740)

the time allowed for completing a certain piece of work is 40 days.
By extending the time by 10 days we find that 25 fewer man will be needed.
How many men are required to finish the work in 40 days?
:
Let x = no. of men required to do the job in 40 days
then we can say
40x man-days required to do the job
:
The same job done in 50 days:
:
50(x-25) = 40x
50x - 1250 = 40x
50x - 40x = 1250
10x = 1250
x = 125 men required to do the job in 40 days
:
:
Check solution by finding the man-days for each crew
40(125) = 5000 md
50(100) = 5000 md